If tana=(1+2^-x)^-1,. tanb=(1+2^(x+1))^-1then. A+b=

Grade:11

2 Answers

Ayanava Datta
6 years ago
answer of the question is 45
hint=find tan(a+b) 
put the value tana ‘tanb
and slove
got tan[a+b]=1
[a+b]=1
 
Arpit Sahu
5 years ago
Tan(a+b)=tana + tanb/1-tanatanb
 =(1+2^-x)^-1 + (1+2^x+1)^-1/1-(1+2^-x)^-1.(1+2^x+1)^-1
 =1+2^x+1+2^-x-1 / 1-(1+2^x)(1+2^-x-1)
=[(2^x+2 + 2^2x+1 + 1)/2^x+1]/[(2^2x+1 + 2^x+1 +1)/2^x+1]
=1
=45°
=π/4